1) A man sells his car for $749 at a loss of 25/2% of what he paid for it. Find his actual loss

2) An article which costs $3.20 in 1940 now costs $9.12. Find the percentage increase in cost.

3) If a dealer charges $2.75 for a chair, he gains 75/2%, at what price must he sell it to gain 45%?


1
Expert's answer
2022-01-18T16:23:28-0500

1)


"100\%-\dfrac{25}{2}\%=87.5"

"\dfrac{749}{87.5\%}\cdot 12.5\%=107"

His actual loss is $107.


2)


"\dfrac{9.12}{3.20}\cdot100\%-100\%=185\%"

The percentage increase in cost is 185%.


3)


"\$2.75\cdot\dfrac{100\%}{\dfrac{75}{2}\%+100\%}=\$2"

"\$2\cdot(1+0.45)=\$2.90"

A dealer must charge $2.90 for a chair.